Hi, I have a question that I really need help with,
The heights of two similar triangles are in the ratio of 2:5. If the area of the larger triangle is 60, then what is the area of the smaller triangle?
I really need help, and would really appreciate it!

Answer by Alan3354(69443)

The heights of two similar triangles are in the ratio of 2:5. If the area of the larger triangle is 60, then what is the area of the smaller triangle?
I really need help, and would really appreciate it!
-----------
The 2 bases are in the same ratio (that's what similar means).
---
Area = b*h/2
For the larger, b*h/2 = 60
For the smaller, it's (2/5)b*(2/5)*h/2
= (4/25)*60
= 9.6 sq units